# Stable Audio Open

### v1

@@ 4.1 Autoencoder training
We train the variational autoencoder using a variety of objectives. First, there is a reconstruction loss, based on a perceptually weighted multi-resolution STFT [23] that deals with stereo signals via the mid-side (M/S) representation of the stereo audio, as well as the left and right (L/R) channels. The L/R component is down-weighted by 0.5 compared to the M/S component, and exists to mitigate ambiguity around left-right placement. Second, we employ an adversarial loss term, utilizing 5 convolutional discriminators as in Encodec [22]. Third, the KL divergence loss term is down-weighted by 1×$\times 10^{-4}$.
@@ 4.2 DiT training and inference
The DiT is trained to predict a noise increment from noised ground-truth latents, following the v-objective [24] approach. During inference, we use the DPM-Solver++ [25] for 100 steps with classifier-free guidance (scale of 0.7). The DiT is trained for 338 h on $\times$64 A100s with a batch size of 4 and a base learning rate of $\times 10^{-5}$. Each batch contains latent sequences of length 1024 ($\approx$47 sec).

### v2

@@ 4.1 Autoencoder training
The (variational) autoencoder is trained using a variety of objectives. First, with a reconstruction loss, based on a perceptually weighted multi-resolution STFT [23] dealing with stereo audio via the mid-side (M/S) and left-right channels (L/R) representations. The L/R component is down-weighted by 0.5 compared to the M/S component, and exists to mitigate ambiguity around left-right placement. Second, we employ an adversarial loss term with feature matching, utilizing 5 convolutional discriminators as in Encodec [22]. Third, we use a KL divergence loss term that is down-weighted by 1×$\times 10^{-4}$.
@@ 4.2 DiT training and inference
The DiT is trained to predict a noise increment from noised ground-truth latents, following the v-objective [24] approach. During inference, we use the DPM-Solver++ [25] for 100 steps with classifier-free guidance (scale of 7.0). The DiT is trained for 338 h on $\times$64 A100s with a batch size of 4 and a base learning rate of $\times 10^{-5}$. Each batch contains latent sequences of length 1024 ($\approx$47 sec).